OpenStudy (vishweshshrimali5):

CHALLENGE QUESTION 3 Let S be a subset of {1,2,3,...,2013} such that no two members of S differ by 4 or 7. The largest number of members of S is divisible by (a) 2 (b) 3 (c) 5 (d) 7

OpenStudy (vishweshshrimali5):

@mukushla @waterineyes

OpenStudy (vishweshshrimali5):

@UnkleRhaukus

OpenStudy (unklerhaukus):

2013/3=671?

OpenStudy (vishweshshrimali5):

yes

OpenStudy (anonymous):

Difference should not be of 4 or 7 @UnkleRhaukus

OpenStudy (unklerhaukus):

huh?

ganeshie8 (ganeshie8):

S can be written as : { 11k + 1, 11k + 2, 11k + 3, 11k + 4, ..... } k = 0, 1, 2, ... .

OpenStudy (anonymous):

Find the multiple of 11 near 2013..

OpenStudy (anonymous):

I think it is the multiple..

OpenStudy (anonymous):

2006 will be the largest number in my opinion..

OpenStudy (anonymous):

And it is divisible by 2 only..

OpenStudy (anonymous):

If I am right then I got the subset S as: S : {1, 2, 3, 4, 12, 13, 14, 15, 23, 24, 25, 26, 34, 35, 36, 37, 45, 46, 47, 48 ......} Now what are the numbers that are missing in it: 5-11 16-22 27-33 38-44 Look for the last pattern: The numbers are divisible by 11.. As 2013 is also divisible by 11 so: The last numbers that will not include in the set are: 2007-2013 So the largest number will be 2006.. May be I am wrong also..

ganeshie8 (ganeshie8):

\(11k + (1 \ to\ 4) \) k = 0 : 1, 2, 3, 4 k = 1 : 12, 13, 14, 15 . . . so from (1 to 2013), for every 11 numbers, we have 4 consecutive numbers that belong to set S so there will be total of (2013/11) * 4 = 732 elements in set S. # numbers divisible by 2 in set S if u have 4 consecutive numbers, only 2 numbers will be divisible by 2. so number of elements in S divisible by 2 = 732/2 = 366

OpenStudy (anonymous):

Question is asking about the largest number..

OpenStudy (anonymous):

So we have to find the largest number of subset S..

OpenStudy (anonymous):

Put k = 183..

ganeshie8 (ganeshie8):

oohhh lol i see it now :) thanks @waterineyes .... i interpreted it in a complicated way

OpenStudy (anonymous):

Sorry not 183.. put k = 182

ganeshie8 (ganeshie8):

yes k = 0 to 182 covers entire set...

OpenStudy (anonymous):

You will get: 2003, 2004, 2005, 2006 So 2006 will be the greatest here..

OpenStudy (anonymous):

For k = 183, 2014, 2015 etc which are not our requirements.. So, 2006 will be the largest number..

OpenStudy (vishweshshrimali5):

@waterineyes kudos to you.......... very gud approach of problem solving.......well regarding the answer and the correct way I will tell you that later...........

OpenStudy (anonymous):

Sure.. This is what Mathematics is.. We can solve the question in many ways.. That is what I Love about Mathematics..

OpenStudy (anonymous):

GOOD JOB!

OpenStudy (anonymous):

@nitz , Using properties of Factorials : \(GOOD \; JOB! = (GOOD \;JOB) \times (GOOD \;JOB - 1)\times(GOOD \;JOB -2).... 1\)

OpenStudy (anonymous):

Hey @vishweshshrimali5 I was just kidding.. Hope you don't mind..

OpenStudy (vishweshshrimali5):

No Problem............ I never mind such things...... Even I liked it much didn't have more medals to give otherwise it would have been yours more........

OpenStudy (vishweshshrimali5):

@waterineyes are you ready for another challenge ?

OpenStudy (anonymous):

I can try for that challenge @vishweshshrimali5 ..